a 5000 kg truck runs into 1000 kg car that was stationary, The truck and car lock together after the collision and move with speed of 8m/s. compute how much kinetic energy was lost in the inelastic collision
1
Expert's answer
2019-11-11T15:54:34-0500

From the conservation of momentum:


"MV=(m+M)v\to (5000)V=(1000+5000)(8)"

"V=9.6\frac{m}{s}"

The kinetic energy was lost in the inelastic collision:


"\Delta E=K_i-K_f=0.5MV^2-0.5(m+M)v^2"

"\Delta E=0.5(5000(9.6)^2-(1000+5000)8^2)=38400\ J"